OpenStudy (anonymous):

For a certain credit card, the total amount, A ,that a customer owes is given by the following formula. In the formula, P is the starting balance and n is the number of months of debt, assuming interest is compounded monthly. Suppose you begin with a credit card debt of $1000. After one year, you will owe $ _______.

OpenStudy (amistre64):

whats the interest rate?

OpenStudy (amistre64):

A = P(1+r/n)^(nt); where n-12 and t = 1 and r = ...something that is not stated yet; and P = 1000 A = 1000(1+r/12)^(12)

OpenStudy (anonymous):

It doesn't say what the interest rate is...but thinks :)

OpenStudy (amistre64):

youre welcome, when it decided to give you an interest rate; all you gotta do is plug it into r :)